please help me simplify this completely. cos B • cot B • sec^2 B
Answer by richwmiller(17219)

cos B • cot B • sec^2 B
adj/hyp*adj/opp *hyp/adj*hyp/adj
a/h*a/o*h/a*h/a=
h/o=
csc(B)